# SoREX: Towards Self-Explainable Social Recommendation with Relevant Ego-Path Extraction

### v1

@@ 4.4. Explanation Re-Aggregation
Explanation re-aggregation aims to aggregate information from sampled ego-paths into original GNN-encoded representations of target users in corresponding tower, such that the factor-specific and candidate-aware knowledge within $\mathcal{G}_{ego}$ can be emphasized. A hop-wise attention based node-level aggregation method is designed for this procedure. We take interaction tower as an example for illustration. Given GNN-encoded user embedding $\mathbf{h}_{i}^{r}$, candidate item $v_{j}$ and interaction-specific $v_{j}$-aware ego-path subset $\tilde{\mathcal{W}}_{ego}^{r}(j)$, we first compute attention in the fashion of Transformer (Vaswani et al., 2017) for each node $q$ in $w_{t}\in\tilde{\mathcal{W}}_{ego}^{r}(j)$ based on its cosine similarity with the GNN-encoded representation of $v_{j}$:
@@ 4.5. Multi-Task Training
where $v$ and $v^{-}$ represent positive and negative sample respectively, and $\mathcal{E}_{t}$ is the training set. The other auxiliary task is friend recommendation task. The introduction of friend recommendation can help distill more supervision signals from social domain in the view of user-user relationships, instead of the relationships between social proximity and user-item interactions only. Recent work (Hu et al., 2023) has demonstrated its positive effect for social recommendation. Therefore, we randomly sample positive and negative user pairs from $\mathcal{G}_{s}$, and conduct candidate user ranking and item ranking simultaneously. Considering that friend recommendation only leverages social networks, we only integrate this task into social tower. The loss function of friend recommendation $\mathcal{L}_{s}$ is defined as:
@@ 4.6. Computational Complexity
This subsection aims to analyze the time and memory complexity of SoREX. Considering that representative baseline methods (e.g. LightGCN (He et al., 2020), Diffnet (Wu et al., 2019) and DESIGN (Tao et al., 2022)) essentially adopt different graph convolution strategies over the same joint graph $\mathcal{G}$, we compare our SoREX with the more computationally efficient method LightGCN and the multi-tower alike method DESIGN. The time and memory complexity of SoREX and the selected baselines are listed in Table 2. Suppose that $|\mathcal{U}|$ and $|\mathcal{V}|$ are the size of user set and item set, while $|\mathcal{E}|$ is the size of edge set of $\mathcal{G}$. Given the dimension of learnable embeddings $d$ and the layer number $L$, the time complexity of LightGCN and SoREX is $O(L|\mathcal{E}|d)$ and $O(L|\mathcal{E}|d+n_{w}|\mathcal{E}_{r}|d)$ respectively. DESIGN essentially performs message propagation over $\mathcal{G}$ for twice, so the time complexity of DESIGN is also $O(L|\mathcal{E}|d)$. The optimal $n_{w}$ usually satisfies $L\ll n_{w}\ll|\mathcal{E}|$. Therefore, the time complexity of SoREX is one to two orders of magnitude greater than that of LightGCN and DESIGN, which is acceptable when dealing with sparse networks. On the other hand, SoREX has no other trainable parameters except for the ID embeddings, which is same as LightGCN and DESIGN. Considering that the ego-path sampling procedures are real-time, the memory costs of SoREX is nearly the twice of LightGCN. Therefore, the memory complexity of SoREX and LightGCN are both $O(L(|\mathcal{U}|+|\mathcal{V}|)d)$. Items are not assigned with ID embeddings in DESIGN, so the memory complexity of DESIGN is $O(L|\mathcal{U}|d)$. In summary, although the time and memory complexity of SoREX are not superior, we aim to achieve the necessary trade-off between efficiency and explainability, while minimizing the unnecessary computational costs.
@@ 5.1. Experimental Setup
Datasets. We evaluate our SoREX on several widely used benchmark datasets, including three sparse datasets Yelp [^1] https://github.com/librahu/HIN-Datasets-for-Recommendation-and-Network-Embedding, Flickr [^2] https://www.flickr.com/, Ciao [^3] https://www.cse.msu.edu/tangjili/datasetcode/truststudy.htm and a dense dataset LastFM [^4] https://files.grouplens.org/datasets/hetrec2011. Yelp is an online location-based social network. Flickr is an online image-based social sharing platform with a whom-trust-whom social network. Ciao is a popular social networking website. LastFM consists of a user-artist interaction network and a user friendship network. Consistent with previous work (Wu et al., 2019), we remove users and items with less than two interaction records. For datasets with ratings like Ciao, we only keep links with ratings no less than 4. Considering that all the selected datasets have no temporal information, we randomly split each dataset into train/validation/test sets at a ratio of 80%/10%/10%. Detailed preprocessed dataset statistics are presented in Table 3.
@@
  • LightGCN (He et al., 2020) is a popular GNN model for general recommendation task. It is characterized by recursive graph convolution without linear transformation and non-linear activation function.
  • LightGCNS is the base model of CGCL (Hu et al., 2023), which is a friend recommendation enhanced LightGCN, conducting additional message propagation on social network to predict social links.
  • SocialLGN (Liao et al., 2022) performs message propagation on both $\mathcal{G}_{r}$ and $\mathcal{G}_{s}$ in each GNN layer to fuse their knowledge.
  • Diffnet (Wu et al., 2019) is a layer-wise social influence propagation model proposed to simulate the social influence diffusion process.
  • Diffnet++ (Wu et al., 2020) is an extension of Diffnet, modeling both social influence diffusion and user preference diffusion in a unified framework.
  • MHCN (Yu et al., 2021b) is a multi-channel hypergraph convolutional network. We remove the self-supervised learning related components in MHCN and test the effectiveness of its proposed model architecture.
  • DESIGN (Tao et al., 2022) leverages multiple GNN encoders to learn from different graphs and fuse their knowledge via knowledge distillation.
  • GBSR (Yang et al., 2024) leverages graph information bottleneck theory to identify and remove redundant social links, so that minimal but sufficient social network could be used for enhanced social recommendation.
@@
  • ConPI (Wang et al., 2021) compares the similarity between neighbor sets of node pairs for both link prediction and explanation. Two versions of ConPI are devised. ConPI-Node identifies similar neighbors as explanations, while ConPI-Pair leverages similar neighbor pairs. We only test ConPI-Node here for the excessive memory complexity of ConPI-Pair.
@@
  • ECF (Du et al., 2023) is a self-explainable collaborative filter method, which aims to discover taste clusters from user-item interactions. The approximation relationships between learned taste clusters and items are used for explanation. Item tags are not adopted for cluster description in our experiment, while only user-item interaction links are used in the evaluation of ECF.
@@
  • GSAT (Miao et al., 2022a) is a self-explainable attention-based GNN proposed to inject learnable randomness into GNN with information bottleneck theory and sample dense explanatory subgraphs without spurious correlations.
@@
  • PxGNN (Dai and Wang, 2024) is a prototype-based self-explainable GNN, which aims to generate prototype graphs for both prediction and prototype-based explanations. To transplant PxGNN into link prediction task, we set the amount of node class as 1, so that all the nodes share the same set of prototype graphs.
@@
Note that both ConPI and GSAT are trained on the joint graph $\mathcal{G}$. We do not adopt other representative self-explainable GNN baselines like SE-GNN (Dai and Wang, 2021), ProtGNN (Zhang et al., 2022), PGIB (Miao et al., 2022b) and DisC (Fan et al., 2024) for comparison, because they are specifically designed for classification tasks. It is infeasible to extract node pair aware explanatory subgraphs via these methods, and thus it is non-trivial to directly transfer them to our task. We also exclude social recommendation baselines SEPT (Yu et al., 2021a), DcRec (Wu et al., 2022) and CGCL (Hu et al., 2023), which focus on self-supervised learning (SSL) for social recommenders. For recent social network denoising based methods (Yang et al., 2024; Li et al., 2024b; Li and Wang, 2024), we only select GBSR for evaluation because their underlying ideas are similar. We would like to compare the superiority of different methods from the perspective of model architecture, so we exclude the SSL related components in the three baselines and test their underlying models, which correspond to LightGCN and LightGCNS.
@@
Evaluation Metrics. Following  (Tao et al., 2022), we adopt Hit Rate (HR@$K$) and Normalized Discounted Cumulative Gain (NDCG@$K$) as evaluation metrics. HR@$K$ evaluates the accuracy of recommendation, while NDCG@$K$ can reflect the quality of item ranking. We set $K=10$ for both HR@$K$ and NDCG@$K$.
@@
Implementation Details. We implement our framework and all selected baselines with PyTorch-Geometric (Fey and Lenssen, 2019). We fix the ID embedding dimension size $d$ as 64 for all methods. For SoREX, hop number of ego-paths $k$, layer number of GNN encoders $k_{1}$ and $k_{2}$ are all tuned from $\{1,2,3\}$. $n_{w}$, the size of $\hat{\mathcal{W}}_{ego}$, is tuned from $\{50,100,150,200,300,500\}$. The friend recommendation task coefficient $\gamma$ is tuned from $\{0.1,0.2,...,1.0\}$. For LightGCN and LightGCNS, we adopt 2-layer GNN architecture. The coefficient of auxiliary task in LightGCNS is tuned from the same range as $\gamma$. For all other baselines, we adopt the recommended settings in their paper. Adam optimizer (Kingma and Ba, 2014) is adopted for the optimization of all models with learning rate 0.001. The batch size is tuned from $\{256,512,1024\}$, and L2 regularization coefficient $\lambda$ is set as 0.001. Following  (Wu et al., 2019), for all methods, we randomly sample 10 and 1000 negative samples for each training and validating sample respectively. During testing, we rank all items that target users have not interacted with. Our code and data are available at https://github.com/antman9914/SoREX.
@@ 5.4. Ablation Study
To demonstrate the effectiveness of our proposed components, we conduct two series of ablation experiments on Yelp dataset: (1) General ablation studies: We first remove auxiliary objective function $\mathcal{L}_{s}$ (w/o $\mathcal{L}_{s}$), our proposed explanation re-aggregation procedure (w/o re-aggr) and all components related to social tower (w/o soc tower) respectively. Then, we replace the social influence aware LightGCN encoder in social tower with the original LightGCN (w/o soc impact). (2) Alternative design studies: We first attempt to replace the ego-path sampling procedure with top-$K$ ego-path extraction to generate sparse explanation graph (w/ top-$K$ ego-path). Then, referring to  (Miao et al., 2022a), we test the effect of additional information bottleneck (IB) based constraints on the sampling probability distributions in both towers, trying to remove potential spurious correlations in explanatory subgraphs (w/ IB constraint). We also try to share the same ID embeddings between the two towers (w/ shared ID), and replace user-transformed item representations $\tilde{h}_{j}^{s}$ used for sampling probability computation in social tower with their corresponding ID embeddings $e_{j}^{s}$ (w/o trans item emb).
@@ 5.6. Explanation Analysis
We analyze our generated explanations from both qualitative and quantitative aspects. We first conduct case study for instance-level qualitative analysis to further illustrate the idea of comparative explanation, and then provide systematic pattern analysis to demonstrate SoREX’s ability to investigate important substructures for recommendation. We also conduct quantitative analysis based on fidelity score (Yeh et al., 2019). All discussions are based on the setting $k=2$. Only testing samples with ground truth items ranked within top-5 are considered.
@@
Quantitative Analysis. We adopt fidelity scores to quantitatively demonstrate the effectiveness of our explanations. Fidelity score measures the performance drop when important features are removed. In our situation, if the model can capture more important ego-paths for each user-item pair, the removal of such ego-paths would lead to more significant performance drop. Considering that GSAT (Miao et al., 2022a) only uses sampled explanatory subgraph for prediction just like our SoREX, we compare SoREX with GSAT and define the fidelity score based on NDCG score. Given user-item pair $(u_{i},v_{j})$, we first re-aggregate the ego-path subset $\hat{\mathcal{W}}_{ego}^{*}\backslash\tilde{\mathcal{W}}_{ego}^{*}(j)$ that are not sampled in each tower to obtain new predicted result $\hat{NDCG}_{ij}$. Then, we calculate the percentage of its performance drop compared with original NDCG@10 result $NDCG_{ij}$. Therefore, the fidelity score $\Delta\mathit{NDCG}\%$ is formulated as $\Delta\mathit{NDCG}\%=\frac{1}{|\mathcal{E}_{t}|}\sum_{(u_{i},v_{j})\in\mathcal{E}_{t}}\frac{(\mathit{NDCG}_{ij}-\hat{\mathit{NDCG}_{ij}})}{NDCG_{ij}}\%$. Similar procedures are also conducted for GSAT. The results on all four datasets are listed in Table 6. Compared with GSAT, SoREX can learn factor-specific and candidate-aware ego-path similarity distributions, while GSAT assigns the same weight for each node regardless of candidate items. Meanwhile, without the re-aggregation module (w/o re-aggr), the model’s fidelity scores decrease significantly, highlighting the critical role of the re-aggregation module in aligning interpretability with prediction results. In addition, the top-$K$ and IB constraint variants show further drops in fidelity, due to their hard truncation and information compression, respectively, both of which lead to loss of important explanatory signals. Thus, our SoREX can provide more appropriate and personalized explanations and consistently achieve higher fidelity scores.

### v2

@@ 4.4. Explanation Re-Aggregation
Explanation re-aggregation aims to aggregate information from sampled ego-paths into original GNN-encoded representations of target users in corresponding tower, such that the factor-specific and candidate-aware knowledge within $\mathcal{G}_{ego}$ can be emphasized. A hop-wise attention based node-level aggregation method is designed for this procedure. We take interaction tower as an example for illustration. Given GNN-encoded user embedding $\mathbf{h}_{i}^{r}$, candidate item $v_{j}$ and interaction-specific $v_{j}$-aware ego-path subset $\tilde{\mathcal{W}}_{ego}^{r}(j)$, we first compute attention in the fashion of Transformer (transformer) for each node $q$ in $w_{t}\in\tilde{\mathcal{W}}_{ego}^{r}(j)$ based on its cosine similarity with the GNN-encoded representation of $v_{j}$:
@@ 4.5. Multi-Task Training
where $v$ and $v^{-}$ represent positive and negative sample respectively, and $\mathcal{E}_{t}$ is the training set. The other auxiliary task is friend recommendation task. The introduction of friend recommendation can help distill more supervision signals from social domain in the view of user-user relationships, instead of the relationships between social proximity and user-item interactions only. Recent work (cgcl) has demonstrated its positive effect for social recommendation. Therefore, we randomly sample positive and negative user pairs from $\mathcal{G}_{s}$, and conduct candidate user ranking and item ranking simultaneously. Considering that friend recommendation only leverages social networks, we only integrate this task into social tower. The loss function of friend recommendation $\mathcal{L}_{s}$ is defined as:
@@ 4.6. Computational Complexity
This subsection aims to analyze the time and memory complexity of SoREX. Considering that representative baseline methods (e.g. LightGCN (lightgcn), Diffnet (diffnet) and DESIGN (design)) essentially adopt different graph convolution strategies over the same joint graph $\mathcal{G}$, we compare our SoREX with the more computationally efficient method LightGCN and the multi-tower alike method DESIGN. The time and memory complexity of SoREX and the selected baselines are listed in Table 2. Suppose that $|\mathcal{U}|$ and $|\mathcal{V}|$ are the size of user set and item set, while $|\mathcal{E}|$ is the size of edge set of $\mathcal{G}$. Given the dimension of learnable embeddings $d$ and the layer number $L$, the time complexity of LightGCN and SoREX is $O(L|\mathcal{E}|d)$ and $O(L|\mathcal{E}|d+n_{w}|\mathcal{E}_{r}|d)$ respectively. DESIGN essentially performs message propagation over $\mathcal{G}$ for twice, so the time complexity of DESIGN is also $O(L|\mathcal{E}|d)$. The optimal $n_{w}$ usually satisfies $L\ll n_{w}\ll|\mathcal{E}|$. Therefore, the time complexity of SoREX is one to two orders of magnitude greater than that of LightGCN and DESIGN, which is acceptable when dealing with sparse networks. On the other hand, SoREX has no other trainable parameters except for the ID embeddings, which is same as LightGCN and DESIGN. Considering that the ego-path sampling procedures are real-time, the memory costs of SoREX is nearly the twice of LightGCN. Therefore, the memory complexity of SoREX and LightGCN are both $O(L(|\mathcal{U}|+|\mathcal{V}|)d)$. Items are not assigned with ID embeddings in DESIGN, so the memory complexity of DESIGN is $O(L|\mathcal{U}|d)$. In summary, although the time and memory complexity of SoREX are not superior, we aim to achieve the necessary trade-off between efficiency and explainability, while minimizing the unnecessary computational costs.
@@ 5.1. Experimental Setup
Datasets. We evaluate our SoREX on several widely used benchmark datasets, including three sparse datasets Yelp [^1] https://github.com/librahu/HIN-Datasets-for-Recommendation-and-Network-Embedding, Flickr [^2] https://www.flickr.com/, Ciao [^3] https://www.cse.msu.edu/tangjili/datasetcode/truststudy.htm and a dense dataset LastFM [^4] https://files.grouplens.org/datasets/hetrec2011. Yelp is an online location-based social network. Flickr is an online image-based social sharing platform with a whom-trust-whom social network. Ciao is a popular social networking website. LastFM consists of a user-artist interaction network and a user friendship network. Consistent with previous work (diffnet), we remove users and items with less than two interaction records. For datasets with ratings like Ciao, we only keep links with ratings no less than 4. Considering that all the selected datasets have no temporal information, we randomly split each dataset into train/validation/test sets at a ratio of 80%/10%/10%. Detailed preprocessed dataset statistics are presented in Table 3.
@@
  • LightGCN (lightgcn) is a popular GNN model for general recommendation task. It is characterized by recursive graph convolution without linear transformation and non-linear activation function.
  • LightGCNS is the base model of CGCL (cgcl), which is a friend recommendation enhanced LightGCN, conducting additional message propagation on social network to predict social links.
  • SocialLGN (sociallgn) performs message propagation on both $\mathcal{G}_{r}$ and $\mathcal{G}_{s}$ in each GNN layer to fuse their knowledge.
  • Diffnet (diffnet) is a layer-wise social influence propagation model proposed to simulate the social influence diffusion process.
  • Diffnet++ (diffnet++) is an extension of Diffnet, modeling both social influence diffusion and user preference diffusion in a unified framework.
  • MHCN (mhcn) is a multi-channel hypergraph convolutional network. We remove the self-supervised learning related components in MHCN and test the effectiveness of its proposed model architecture.
  • DESIGN (design) leverages multiple GNN encoders to learn from different graphs and fuse their knowledge via knowledge distillation.
  • GBSR (gbsr) leverages graph information bottleneck theory to identify and remove redundant social links, so that minimal but sufficient social network could be used for enhanced social recommendation.
@@
  • ConPI (conpi) compares the similarity between neighbor sets of node pairs for both link prediction and explanation. Two versions of ConPI are devised. ConPI-Node identifies similar neighbors as explanations, while ConPI-Pair leverages similar neighbor pairs. We only test ConPI-Node here for the excessive memory complexity of ConPI-Pair.
@@
  • ECF (ecf) is a self-explainable collaborative filter method, which aims to discover taste clusters from user-item interactions. The approximation relationships between learned taste clusters and items are used for explanation. Item tags are not adopted for cluster description in our experiment, while only user-item interaction links are used in the evaluation of ECF.
@@
  • GSAT (gsat) is a self-explainable attention-based GNN proposed to inject learnable randomness into GNN with information bottleneck theory and sample dense explanatory subgraphs without spurious correlations.
@@
  • PxGNN (pxgnn) is a prototype-based self-explainable GNN, which aims to generate prototype graphs for both prediction and prototype-based explanations. To transplant PxGNN into link prediction task, we set the amount of node class as 1, so that all the nodes share the same set of prototype graphs.
@@
Note that both ConPI and GSAT are trained on the joint graph $\mathcal{G}$. We do not adopt other representative self-explainable GNN baselines like SE-GNN (se-gnn), ProtGNN (protgnn), PGIB (lri) and DisC (disc) for comparison, because they are specifically designed for classification tasks. It is infeasible to extract node pair aware explanatory subgraphs via these methods, and thus it is non-trivial to directly transfer them to our task. We also exclude social recommendation baselines SEPT (sept), DcRec (derec) and CGCL (cgcl), which focus on self-supervised learning (SSL) for social recommenders. For recent social network denoising based methods (gbsr; recdiff; gdssl), we only select GBSR for evaluation because their underlying ideas are similar. We would like to compare the superiority of different methods from the perspective of model architecture, so we exclude the SSL related components in the three baselines and test their underlying models, which correspond to LightGCN and LightGCNS.
@@
Evaluation Metrics. Following  (design), we adopt Hit Rate (HR@$K$) and Normalized Discounted Cumulative Gain (NDCG@$K$) as evaluation metrics. HR@$K$ evaluates the accuracy of recommendation, while NDCG@$K$ can reflect the quality of item ranking. We set $K=10$ for both HR@$K$ and NDCG@$K$.
@@
Implementation Details. We implement our framework and all selected baselines with PyTorch-Geometric (pyg_lib). We fix the ID embedding dimension size $d$ as 64 for all methods. For SoREX, hop number of ego-paths $k$, layer number of GNN encoders $k_{1}$ and $k_{2}$ are all tuned from $\{1,2,3\}$. $n_{w}$, the size of $\hat{\mathcal{W}}_{ego}$, is tuned from $\{50,100,150,200,300,500\}$. The friend recommendation task coefficient $\gamma$ is tuned from $\{0.1,0.2,...,1.0\}$. For LightGCN and LightGCNS, we adopt 2-layer GNN architecture. The coefficient of auxiliary task in LightGCNS is tuned from the same range as $\gamma$. For all other baselines, we adopt the recommended settings in their paper. Adam optimizer (adam) is adopted for the optimization of all models with learning rate 0.001. The batch size is tuned from $\{256,512,1024\}$, and L2 regularization coefficient $\lambda$ is set as 0.001. Following  (diffnet), for all methods, we randomly sample 10 and 1000 negative samples for each training and validating sample respectively. During testing, we rank all items that target users have not interacted with. Our code and data are available at https://github.com/antman9914/SoREX.
@@ 5.4. Ablation Study
To demonstrate the effectiveness of our proposed components, we conduct two series of ablation experiments on Yelp dataset: (1) General ablation studies: We first remove auxiliary objective function $\mathcal{L}_{s}$ (w/o $\mathcal{L}_{s}$), our proposed explanation re-aggregation procedure (w/o re-aggr) and all components related to social tower (w/o soc tower) respectively. Then, we replace the social influence aware LightGCN encoder in social tower with the original LightGCN (w/o soc impact). (2) Alternative design studies: We first attempt to replace the ego-path sampling procedure with top-$K$ ego-path extraction to generate sparse explanation graph (w/ top-$K$ ego-path). Then, referring to  (gsat), we test the effect of additional information bottleneck (IB) based constraints on the sampling probability distributions in both towers, trying to remove potential spurious correlations in explanatory subgraphs (w/ IB constraint). We also try to share the same ID embeddings between the two towers (w/ shared ID), and replace user-transformed item representations $\tilde{h}_{j}^{s}$ used for sampling probability computation in social tower with their corresponding ID embeddings $e_{j}^{s}$ (w/o trans item emb).
@@ 5.6. Explanation Analysis
We analyze our generated explanations from both qualitative and quantitative aspects. We first conduct case study for instance-level qualitative analysis to further illustrate the idea of comparative explanation, and then provide systematic pattern analysis to demonstrate SoREX’s ability to investigate important substructures for recommendation. We also conduct quantitative analysis based on fidelity score (fidelity). All discussions are based on the setting $k=2$. Only testing samples with ground truth items ranked within top-5 are considered.
@@
Quantitative Analysis. We adopt fidelity scores to quantitatively demonstrate the effectiveness of our explanations. Fidelity score measures the performance drop when important features are removed. In our situation, if the model can capture more important ego-paths for each user-item pair, the removal of such ego-paths would lead to more significant performance drop. Considering that GSAT (gsat) only uses sampled explanatory subgraph for prediction just like our SoREX, we compare SoREX with GSAT and define the fidelity score based on NDCG score. Given user-item pair $(u_{i},v_{j})$, we first re-aggregate the ego-path subset $\hat{\mathcal{W}}_{ego}^{*}\backslash\tilde{\mathcal{W}}_{ego}^{*}(j)$ that are not sampled in each tower to obtain new predicted result $\hat{NDCG}_{ij}$. Then, we calculate the percentage of its performance drop compared with original NDCG@10 result $NDCG_{ij}$. Therefore, the fidelity score $\Delta\mathit{NDCG}\%$ is formulated as $\Delta\mathit{NDCG}\%=\frac{1}{|\mathcal{E}_{t}|}\sum_{(u_{i},v_{j})\in\mathcal{E}_{t}}\frac{(\mathit{NDCG}_{ij}-\hat{\mathit{NDCG}_{ij}})}{NDCG_{ij}}\%$. Similar procedures are also conducted for GSAT. The results on all four datasets are listed in Table 6. Compared with GSAT, SoREX can learn factor-specific and candidate-aware ego-path similarity distributions, while GSAT assigns the same weight for each node regardless of candidate items. Meanwhile, without the re-aggregation module (w/o re-aggr), the model’s fidelity scores decrease significantly, highlighting the critical role of the re-aggregation module in aligning interpretability with prediction results. In addition, the top-$K$ and IB constraint variants show further drops in fidelity, due to their hard truncation and information compression, respectively, both of which lead to loss of important explanatory signals. Thus, our SoREX can provide more appropriate and personalized explanations and consistently achieve higher fidelity scores.